if the mth term of an a.p is n and nth term of an a.p is m then (m+n)th term is what?

Grade:10

1 Answers

Raghav Kumar
7 years ago
a+(m-1)d=n{given}a+(n-1)d=m{given}By subtracting above two equation ,we getd=-1&a=m+n-1Now,(m+n)th term will be a+(m+n-1)d=m+n-1-m-n+1=0[by putting the value of a and d]